# Discovering Governing equations from Graph-Structured Data by Sparse Identification of Nonlinear Dynamical Systems

### v1

@@ Abstract
The combination of machine learning (ML) and sparsity-promoting techniques is enabling direct extraction of governing equations from data, revolutionizing computational modeling in diverse fields of science and engineering. The discovered dynamical models could be used to address challenges in climate science, neuroscience, ecology, finance, epidemiology, and beyond. However, most existing sparse identification methods for discovering dynamical systems treat the whole system as one without considering the interactions between subsystems. As a result, such models are not able to capture small changes in the emergent system behavior. To address this issue, we developed a new method called Sparse Identification of Nonlinear Dynamical Systems from Graph-structured data (SINDyG), which incorporates the network structure into sparse regression to identify model parameters that explain the underlying network dynamics. SINDyG discovers the governing equations of network dynamics while offering improvements in accuracy and model simplicity.
Index Terms:
Data-Driven Modeling, Network Dynamics, Sparse Regression
@@ 1 Introduction
Despite the encouraging results by models like SINDy, they are not designed for complex networked systems with many interacting components such as biological neural networks in the brain [10], power grids [11], social networks [12], and traffic systems [13]. In these systems, the state of each node is influenced not only by its own dynamics but also by the interactions with its connected neighbors [14]. This is because complex system behavior arises from the dynamic interactions between the nodes, not just from the nodes themselves [15]. In this paper, we propose an extension of the SINDy method, termed SINDyG (Sparse Identification of Nonlinear Dynamics for Graph-structured data), which explicitly accounts for the network structure during sparse regression. By incorporating the connections between nodes, our approach uncovers not only the individual dynamics of each node but also the mechanisms by which they interact, leading to more accurate and interpretable models. This is particularly important as the dynamics of a state can propagate through connections, influencing the behavior of the entire system [16].
The work is organized as follows: In Sec. II, we discuss the Sparse Identification of Nonlinear Dynamics (SINDy) [7] and all the necessary information about graph-structured data. In Sec. III, we explain our proposed method for data-driven modeling of complex and interconnected dynamical systems. Next, Section IV presents the results of our method using several case studies based on neuronal dynamics. Finally, we conclude the paper in Section V.
@@ 2.1 Sparse Identification of Nonlinear Dynamics (SINDy)
In order to compute the function $F$ from data, a time series of system’s state variables $\mathbf{x}(t)$ is collected. Also, the derivatives of state variables $\mathbf{\dot{x}}(t)$ are measured or approximated numerically from $\mathbf{x}(t)$ and arranged in the following matrix:
@@
Afterwards, a library $\Theta(\mathbf{X})$ consisting of candidate nonlinear functions of the columns of $\mathbf{X}$ is constructed. For instance, $\Theta(\mathbf{X})$ may consist of constant, polynomial, and trigonometric terms:
@@
$\Theta(\mathbf{X})=\left[\begin{array}[]{cccccccc}\mid&\mid&\mid&\mid&&\mid&\mid\\ 1&\mathbf{X}&\mathbf{X}^{P_{2}}&\mathbf{X}^{P_{3}}&\cdots&\sin(\mathbf{X})&\cos(\mathbf{X})&\cdots\\ \mid&\mid&\mid&\mid&&\mid&\mid&\end{array}\right]$
@@
where, higher polynomials are shown as $\mathbf{X}^{P_{2}}$, $\mathbf{X}^{P_{3}}$, $\cdots$. For example, $\mathbf{X}^{P_{2}}$ denotes the quadratic nonlinearities in the state $\mathbf{x}$. A potential function for the right-hand side of Eq 1 is represented by each column of $\Theta(\mathbf{X})$ in Eq 3. The values of each column are often normalized to prevent numerical instability and ensure that features contribute equally to the model fitting process. Since only a small number of these terms are active in each row of $F$, a sparse regression problem will be established to find the sparse vectors of coefficients $\mathbf{\Xi}=\left[\begin{array}[]{llll}\boldsymbol{\xi}_{1}&\boldsymbol{\xi}_{2}&\cdots&\boldsymbol{\xi}_{K}\end{array}\right]$ that determine which nonlinear terms are active:
$\dot{\mathbf{X}}=\Theta(\mathbf{X})\mathbf{\Xi}$ (4)
@@
For each column $\boldsymbol{\xi}_{k}$ of $\mathbf{\Xi}$, there is a sparse vector of coefficients that determine which terms are active in the righthand side for one of the row equations $\dot{\mathbf{x}}_{k}=\mathbf{F}_{k}\left(\mathbf{x}\right)$ in Eq 1. After $\mathbf{\Xi}$ is identified, a model of each row of the governing equations can be built as follows:
$\dot{\mathbf{x}}_{k}=\mathbf{F}_{k}\left(\mathbf{x}\right)=\Theta({\mathbf{x}}^{T})\boldsymbol{\xi}_{k}$ (5)
@@
Unlike $\Theta(\mathbf{X})$, which is a data matrix, $\Theta({\mathbf{x}}^{T})$ is a vector of symbolic functions of elements of ${\mathbf{x}}$. As a result,
$\dot{\mathbf{x}}=\mathbf{F}\left(\mathbf{x}\right)=\mathbf{\Xi}^{T}\left(\Theta({\mathbf{x}}^{T})\right)^{T}$ (6)
@@
To obtain the sparse vector of coefficients $\boldsymbol{\xi}_{k}$ for the kth row equation, each column of $\dot{\mathbf{X}}$ in Eq 4 requires a distinct optimization. Furthermore, the optimization is given by
$\min_{\boldsymbol{\Xi}}\frac{1}{2}\|\dot{\mathbf{X}}-\boldsymbol{\Theta}(\mathbf{X})\boldsymbol{\Xi}\|^{2}+\lambda R(\boldsymbol{\Xi})$ (7)
@@
where $\lambda$ is a hyper-parameter that controls the regularization’s strength and $R(.)$ is a regularizer that encourages sparsity. There are several well-known techniques for Eq 7 when $R(.)$ is convex. The typical technique is to use $R(.)$ to represent the sparsity-promoting $L1$ norm, which is a convex relaxation of the $L0$ norm. While LASSO [17] is commonly used to solve SINDy due to its convex formulation, it has several limitations such as bias in coefficient estimation. Hence, the SINDy algorithm employs sequential thresholded least squares (STLSQ), which given a parameter $\eta$ that specifies the minimum magnitude for a coefficient in $\mathbf{\Xi}$, performs a least squares fit and then zeros out all the coefficients with magnitude below the threshold. This process of fitting and thresholding is repeated until convergence. The SINDy algorithm has recently been integrated into a python package PySINDy [18].
@@
Figure 1: Overview of the proposed framework for network dynamics discovery. Step 1: Data preparation involves collecting time-series data $\mathbf{X}$ from a graph-structured system. Step 2: The SINDyG algorithm constructs a library of basis functions $\boldsymbol{\Theta}(\mathbf{X})$ and employs sparse regression to identify the optimal sparse coefficient matrix $\boldsymbol{\Xi}$ that satisfies $\dot{\mathbf{X}}=\boldsymbol{\Theta}(\mathbf{X})\boldsymbol{\Xi}$, while taking into account the network structure during the process. Step 3: The predicted dynamics are validated using appropriate metrics to assess model performance  [7] [19].
@@ 2.2 Graph Structured Time Series Data
We represent the connections in the graph using an adjacency matrix $A$, where each element $A_{ij}$ denotes the presence of a connection from node $i$ to node $j$. In a directed graph, $A_{ij}$ can differ from $A_{ji}$, indicating that the influence from node $i$ to node $j$ is not necessarily reciprocal. Connections in the graph, represented by the edges, signify the pathways through which information or influence is transmitted between nodes. These connections can be weighted to represent the varying strengths of interactions. For example, in a neural network, a stronger synaptic connection might result in a more significant influence of one neuron’s activity on another.
@@ 3.1 Method
The SINDyG method is based on SINDy algorithm which is described in Section 2. In SINDyG, shown in figure 1, the time history of the state variables $\mathbf{X}$ are measured, their derivatives are calculated, and the structure of the network identified. Next, a library of functions of the states, $\boldsymbol{\Theta}(\mathbf{X})$, is constructed. This feature library is used to find the fewest terms needed to satisfy $\dot{\mathbf{X}}=\boldsymbol{\Theta}(\mathbf{X})\boldsymbol{\Xi}$ while taking into account the network structure. The few entries in the vectors of $\boldsymbol{\Xi}$ are identified using sparse regression [7]. Here, we provide details of each step of the SINDyG method.
@@ 3.1.1 Data Preparation
Data preparation is a crucial step in the discovery of dynamical systems using SINDyG method. This preparation involves gathering and organizing three key components: an adjacency matrix, time series data, and their derivatives. These components collectively provide the necessary information to uncover the underlying equations governing the system’s behavior.
@@ 3.1.2 Network Dynamics Discovery
In this step, we aim to uncover the underlying dynamical equations governing each node in the network. We begin by constructing a library matrix, ${\Theta}({X})$, using a predefined set of basis functions. These basis functions represent potential terms that could contribute to the dynamics [7]. Next, we employ a sequentially thresholded least squares algorithm (STLSQ) [7] to determine the coefficients $\Xi$ corresponding to these basis functions. This is achieved by minimizing the following objective function:
$\min_{\Xi}\frac{1}{2}\|\dot{{X}}-{\Theta}({X})\Xi\|_{2}^{2}+\lambda\|f\cdot\Xi\|_{2}^{2}$ (8)
@@
The parameter $\lambda$ controls the sparsity of the solution, while the function $f$ leverages the network’s adjacency matrix $(A)$ to introduce a graph-aware penalty. This penalty term favors coefficients associated with strongly connected nodes and discourages those associated with weakly connected nodes, guiding the algorithm towards solutions that align with the network’s structure. The intuition behind function $f$ is to determine the relative importance of different terms in the library based on the network’s connectivity. By establishing a mathematical relationship between the adjacency matrix and the desired active terms, we ensure that the discovered dynamics are consistent with the underlying network topology.
@@
In order to calculate this penalty, we iterate through each candidate term, $j$, in the library of functions. For each term, we identify the source indices, $S_{j}$, representing the state variables present in term $j$. We then determine the sink indices, $D_{j}$, which are the state variables reachable from $S_{j}$. For each sink index $i\in D_{j}$, we compute the mean connectivity, $m_{ij}$, between the source group(s) and sink $i$ using the adjacency matrix $A$. The penalty is then calculated using a function similar to the sigmoid function:
$f[i,j]=\frac{1}{1+\exp\left(\left(L/\left|S_{j}\right|\right)\cdot\left(m_{ij}-0.5\right)\right)}$ (9)
@@
where $L$ is a parameter for adjusting the shape of this function. As we increase $L$, the function becomes more similar to a step function. Conversely, if we reduce $L$, the function’s behavior resembles a ramp function. This formulation ensures that terms with low mean connectivity between source and sink variables receive a higher penalty (closer to 1), while the penalty is scaled by the number of source variables, $\left|S_{j}\right|$, to account for terms involving multiple state variables. The shape of the function $f[i,j]$ based on the value of $L/\left|S_{j}\right|$ is illustrated in figure  2.
@@
Figure 2: Comparison of $f[i,j]$ with different values of $L/\left|S_{j}\right|$
@@
To solve the optimization problem shown in Eq. 8 for each state variable $i$, we perform a series of transformations that allow us to leverage the standard ridge regression framework. First, we scale the feature matrix ${\Theta}({X})$ and the coefficient vector $\Xi_{i}$ to simplify the regularization term. Specifically, we define the transformed variables ${\Theta}({X})^{\prime}$ and ${\Xi_{i}}^{\prime}$ as follows: ${\Theta}({X})^{\prime}={\Theta}({X})\cdot\operatorname{diag}(1/f_{i})$, where each column $j$ of ${\Theta}({X})$ is divided by the corresponding element $f_{ij}$, and ${\Xi_{i}}^{\prime}=\operatorname{diag}(f_{i})\cdot\Xi_{i}$, which scales each coefficient in $\Xi_{i}$ by the corresponding element in $f_{i}$. These substitutions transform the original problem into the equivalent form $\|\dot{{X_{i}}}-{\Theta}({X})^{\prime}{\Xi_{i}}^{\prime}\|_{2}^{2}+\lambda\|{\Xi_{i}}^{\prime}\|_{2}^{2}$.
@@
In this transformed space, the problem becomes a standard ridge regression, which can be solved efficiently using existing algorithms. After obtaining the solution ${\Xi_{i}}^{\prime}$ using sequentially thresholded least squares algorithm, we revert to the original coefficient vector $\Xi_{i}$ by dividing ${\Xi_{i}}^{\prime}$ by $f_{i}$, ensuring that the final coefficients correspond to the original problem’s scaling. This approach effectively incorporates the custom regularization term $\lambda\|f_{i}\cdot\Xi_{i}\|_{2}^{2}$ into the ridge regression framework without altering the underlying algorithm. By scaling the feature matrix and coefficients appropriately, we maintain the efficiency of the standard ridge regression solver while addressing the modified regularization objective.

### v2

# M.A. Basiri & S. Khanmohammadi

@@ Abstract
The combination of machine learning (ML) and sparsity-promoting techniques is enabling direct extraction of governing equations from data, revolutionizing computational modeling in diverse fields of science and engineering. The discovered dynamical models could be used to address challenges in climate science, neuroscience, ecology, finance, epidemiology, and beyond. However, most existing sparse identification methods for discovering dynamical systems treat the whole system as one without considering the interactions between subsystems. As a result, such models are not able to capture small changes in the emergent system behavior. To address this issue, we developed a new method called Sparse Identification of Nonlinear Dynamical Systems from Graph-structured data (SINDyG), which incorporates the network structure into sparse regression to identify model parameters that explain the underlying network dynamics. We showcase the application of our proposed method using several case studies of neuronal dynamics, where we model the macroscopic oscillation of a population of neurons using the extended Stuart-Landau (SL) equation and utilize the SINDyG method to identify the underlying nonlinear dynamics. Our extensive computational experiments validate the improved accuracy and simplicity of discovered network dynamics when compared to the original SINDy approach. Data-Driven Modeling, Sparse Regression, Network Dynamics, Coupled Oscillators, Neuronal Population
@@ 1 Introduction
Despite the encouraging results by models like SINDy, they are not designed for complex networked systems with many interacting components such as biological neural networks in the brain [15], power grids [18], social networks [1], and traffic systems [24]. In these systems, the state of each node is influenced not only by its own dynamics but also by the interactions with its connected neighbors [22]. This is because complex system behavior arises from the dynamic interactions between the nodes, not just from the nodes themselves [9]. Several methods have been developed to infer network dynamics while taking into account the interaction between nodes [9, 4]. However, majority of these methods typically do not leverage existing knowledge of graph structures or directly incorporate them into the modeling of node-level dynamics.
In this paper, we propose an extension of the SINDy method, termed SINDyG (Sparse Identification of Nonlinear Dynamics for Graph-structured data), which explicitly accounts for the network structure during sparse regression. By incorporating the connections between nodes, our approach uncovers not only the individual dynamics of each node but also the mechanisms by which they interact, leading to more accurate and interpretable models. We test our method using several case studies of neuronal dynamics, where the goal is to identify nonlinear interactions between populations of neurons. Our results indicate improvements to model accuracy and simplicity when compared to the original SINDy method.
The rest of the work is organized as follows: In Sec. II, we discuss the Sparse Identification of Nonlinear Dynamics (SINDy) [3] and all the necessary information about graph-structured data. In Sec. III, we explain our proposed method for data-driven modeling of complex and interconnected dynamical systems. Next, Section IV presents the results of our method using several case studies based on neuronal dynamics. Finally, we conclude the paper in Section V.
@@ 2.1 Sparse Identification of Nonlinear Dynamics (SINDy)
In order to compute the function $\mathbf{F}$ from data, a time series of system’s state variables $\mathbf{x}(t)$ is collected. Also, the derivatives of state variables $\mathbf{\dot{x}}(t)$ are measured or approximated numerically from $\mathbf{x}(t)$ and arranged in the following matrix:
@@
Afterwards, a library $\mathbf{\Theta}(\mathbf{X})$ consisting of candidate nonlinear functions of the columns of $\mathbf{X}$ is constructed. For instance, $\mathbf{\Theta}(\mathbf{X})$ may consist of constant, polynomial, and trigonometric terms:
@@
$\mathbf{\Theta}(\mathbf{X})=\left[\begin{array}[]{cccccccc}\mid&\mid&\mid&\mid&&\mid&\mid\\ 1&\mathbf{X}&\mathbf{X}^{P_{2}}&\mathbf{X}^{P_{3}}&\cdots&\sin(\mathbf{X})&\cos(\mathbf{X})&\cdots\\ \mid&\mid&\mid&\mid&&\mid&\mid&\end{array}\right]$
@@
where, higher polynomials are shown as $\mathbf{X}^{P_{2}}$, $\mathbf{X}^{P_{3}}$, $\cdots$. For example, $\mathbf{X}^{P_{2}}$ denotes the quadratic nonlinearities in the state $\mathbf{x}$. A potential function for the right-hand side of Eq 1 is represented by each column of $\mathbf{\Theta}(\mathbf{X})$ in Eq 3. The values of each column are often normalized to prevent numerical instability and ensure that features contribute equally to the model fitting process. Since only a small number of these terms are active in each row of $\mathbf{F}$, a sparse regression problem will be established to find the sparse vectors of coefficients $\mathbf{\Xi}=\left[\begin{array}[]{llll}\bm{\xi}_{1}&\bm{\xi}_{2}&\cdots&\bm{\xi}_{K}\end{array}\right]$ that determine which nonlinear terms are active:
$\dot{\mathbf{X}}=\mathbf{\Theta}(\mathbf{X})\mathbf{\Xi}$ (4)
@@
For each column $\bm{\xi}_{k}$ of $\mathbf{\Xi}$, there is a sparse vector of coefficients that determine which terms are active in the righthand side for one of the row equations $\dot{\mathbf{x}}_{k}=\mathbf{F}_{k}\left(\mathbf{x}\right)$ in Eq 1. After $\mathbf{\Xi}$ is identified, a model of each row of the governing equations can be built as follows:
$\dot{\mathbf{x}}_{k}=\mathbf{F}_{k}\left(\mathbf{x}\right)=\mathbf{\Theta}({\mathbf{x}}^{T})\bm{\xi}_{k}$ (5)
@@
Unlike $\mathbf{\Theta}(\mathbf{X})$, which is a data matrix, $\mathbf{\Theta}({\mathbf{x}}^{T})$ is a vector of symbolic functions of elements of ${\mathbf{x}}$. As a result,
$\dot{\mathbf{x}}=\mathbf{F}\left(\mathbf{x}\right)=\mathbf{\Xi}^{T}\left(\mathbf{\Theta}({\mathbf{x}}^{T})\right)^{T}$ (6)
@@
To obtain the sparse vector of coefficients $\bm{\xi}_{k}$ for the kth row equation, each column of $\dot{\mathbf{X}}$ in Eq 4 requires a distinct optimization. Furthermore, the optimization is given by
$\min_{\mathbf{\Xi}}\frac{1}{2}\|\dot{\mathbf{X}}-\mathbf{\Theta}(\mathbf{X})\mathbf{\Xi}\|^{2}+\lambda\mathbf{R}(\mathbf{\Xi})$ (7)
@@
where $\lambda$ is a hyper-parameter that controls the regularization’s strength and $\mathbf{R}(.)$ is a regularizer that imposes sparsity or other constraints on the solution. There are several well-known techniques for Eq 7 when $\mathbf{R}(.)$ is convex. The typical technique is to use $\mathbf{R}(.)$ to represent the sparsity-promoting $L1$ norm, which is a convex relaxation of the $L0$ norm. While LASSO [12] is commonly used to solve SINDy due to its convex formulation, it has several limitations such as bias in coefficient estimation. Hence, the SINDy algorithm employs sequential thresholded least squares (STLSQ), which given a parameter $\eta$ that specifies the minimum magnitude for a coefficient in $\mathbf{\Xi}$, performs a least squares fit and then zeros out all the coefficients with magnitude below the threshold. This process of fitting and thresholding is repeated until convergence. The SINDy algorithm has recently been integrated into a python package PySINDy [6].
@@
Figure 1: Overview of the proposed framework for network dynamics discovery: (Step 1) Data preparation involves collecting time-series data $\mathbf{X}$ from a graph-structured system. (Step 2) The SINDyG algorithm constructs a library of basis functions $\mathbf{\Theta}(\mathbf{X})$, applying sparse regression via the objective function and STLSQ procedure to identify the optimal sparse coefficient matrix $\mathbf{\Xi}$ that satisfies $\dot{\mathbf{X}}=\mathbf{\Theta}(\mathbf{X})\mathbf{\Xi}$ while incorporating the network structure. (Step 3) The predicted dynamics are validated using appropriate metrics to evaluate model performance  [3] [5].
@@ 2.2 Graph Structured Time Series Data
We represent the connections in the graph using an adjacency matrix $\mathbf{A}$, where each element $\mathbf{A}_{mn}$ denotes the presence of a connection from node $m$ to node $n$. In a directed graph, $\mathbf{A}_{mn}$ can differ from $\mathbf{A}_{nm}$, indicating that the influence from node $m$ to node $n$ is not necessarily reciprocal. Connections in the graph, represented by the edges, signify the pathways through which information or influence is transmitted between nodes. These connections can be weighted to represent the varying strengths of interactions. For example, in a neural network, a stronger synaptic connection might result in a more significant influence of one neuron’s activity on another.
@@ 3.1 Method
The SINDyG method is based on the SINDy algorithm (see Section 2). In SINDyG, shown in figure 1, the time histories of the state variables $\mathbf{X}$ are collected, their derivatives are calculated, and the structure of the network is identified. Next, a library of functions of the states, $\mathbf{\Theta}(\mathbf{X})$, is constructed. This feature library is used to find the fewest terms needed to satisfy $\dot{\mathbf{X}}=\mathbf{\Theta}(\mathbf{X})\mathbf{\Xi}$ while taking into account the network structure. The few active entries in the vectors of $\mathbf{\Xi}$ are identified using sparse regression [3]. Here, we provide details of each step of the SINDyG method.
@@ 3.1.1 Data Preparation
Data preparation is a crucial step in the discovery of dynamical systems using the SINDyG method. This preparation involves gathering and organizing three key components: an adjacency matrix, time series data, and their derivatives. These components collectively provide the necessary information to uncover the underlying equations governing the system’s behavior.
@@ 3.1.2 Network Dynamics Discovery
In this step, we aim to uncover the underlying dynamical equations governing each node in the network. We begin by constructing a library matrix, $\mathbf{\Theta}(\mathbf{{X}})$, using a predefined set of basis functions. These basis functions represent potential terms that could contribute to the dynamics [3]. Next, we employ a sequentially thresholded least squares algorithm (STLSQ) [3] to determine the coefficients $\mathbf{\Xi}$ corresponding to these basis functions. This is achieved by minimizing the following objective function:
$\min_{\mathbf{\Xi}}\frac{1}{2}\|\dot{\mathbf{X}}-\mathbf{\Theta}(\mathbf{X})\mathbf{\Xi}\|_{2}^{2}+\lambda\|\mathbf{P}\odot\mathbf{\Xi}\|_{2}^{2}$ (8)
@@
The parameter $\lambda$ controls the amount of regularization. The term $\mathbf{P}\odot\mathbf{\Xi}$ represents the element-wise product of $\mathbf{P}$ and $\mathbf{\Xi}$, while the penalty matrix $\mathbf{P}$ leverages the network’s adjacency matrix $\mathbf{A}$ to introduce a graph-aware penalty. The intuition behind the penalty matrix $\mathbf{P}$ is to determine the relative importance of different terms in the library based on the network’s connectivity. By encouraging coefficients linked to strongly connected nodes and discouraging those linked to weakly connected nodes, this penalty term guides the algorithm toward solutions that ensure the discovered dynamics are consistent with the underlying network topology.
@@
In order to calculate this penalty, we iterate through each candidate term, $c$, in the library of functions. For each term, we identify the source indices, $S_{c}$, representing the state variables present in term $c$. We then determine the sink indices, $D_{c}$, which are the state variables reachable from $S_{c}$. For each sink index $k\in D_{c}$, we compute the mean connectivity, $m_{ck}$, between the source indices and sink index $k$ using the adjacency matrix $\mathbf{A}$. The penalty matrix is then calculated using a formula similar to the sigmoid formula:
$\mathbf{P}[c,k]=\frac{1}{1+\exp\left(\left(L/\left|S_{c}\right|\right)\cdot\left(m_{ck}-0.5\right)\right)}$ (9)
@@
where $L$ is a parameter for adjusting the shape of the output of this formula. As we increase $L$, the behavior of the equation becomes more similar to a step function. Conversely, if we reduce $L$, the equation’s behavior resembles a ramp function. This formulation ensures that terms with low mean connectivity between source and sink variables receive a higher penalty (closer to 1), while the penalty is scaled by the number of source variables, $\left|S_{c}\right|$, to account for terms involving multiple state variables. The shape of the value of $\mathbf{P}[c,k]$ based on the value of $L/\left|S_{c}\right|$ is illustrated in figure 2.
@@
Figure 2: Comparison of $\mathbf{P}[c,k]$ with different values of $L/\left|S_{c}\right|$ and $m_{ck}$
@@
To solve the optimization problem shown in Eq. 8 for each state variable $k$, we perform a series of transformations that allow us to leverage the standard ridge regression framework. First, we scale the feature matrix $\mathbf{\Theta}(\mathbf{X})$ and the coefficient vector $\bm{\xi}_{k}$ to simplify the regularization term. Specifically, we define the transformed variables $\mathbf{\Theta}^{\prime}(\mathbf{X})$ and ${\bm{\xi}_{k}^{\prime}}$ as follows:
$\mathbf{\Theta}^{\prime}(\mathbf{X})=\mathbf{\Theta}(\mathbf{X})\cdot\operatorname{diag}(1/\bm{p}_{k}),\quad{\bm{\xi}_{k}^{\prime}}=\operatorname{diag}(\bm{p}_{k})\cdot\bm{\xi}_{k},$
where each column $c$ of $\mathbf{\Theta}(\mathbf{X})$ is divided by the corresponding element $\bm{P}_{ck}$, and ${\bm{\xi}_{k}^{\prime}}=\operatorname{diag}(\bm{p}_{k})\cdot\bm{\xi}_{k}$, which scales each coefficient in $\bm{\xi}_{k}$ by the corresponding element in $\bm{p}_{k}$ which is column $k$ of $\mathbf{P}$. These substitutions transform the original problem into the equivalent form $\|\dot{\mathbf{x}}_{k}-\mathbf{\Theta}^{\prime}(\mathbf{X}){\bm{\xi}_{k}^{\prime}}\|_{2}^{2}+\lambda\|{\bm{\xi}_{k}^{\prime}}\|_{2}^{2}$.
@@
In this transformed space, the problem becomes a standard ridge regression, which can be solved efficiently using existing algorithms. After obtaining the solution ${\bm{\xi}_{k}^{\prime}}$ using the sequentially thresholded least squares algorithm, we revert to the original coefficient vector $\bm{\xi}_{k}$ by dividing ${\bm{\xi}_{k}^{\prime}}$ by $\bm{p}_{k}$, ensuring that the final coefficients correspond to the original problem’s scaling. This approach effectively incorporates the custom regularization term $\lambda\|\bm{p}_{k}\odot\bm{\xi}_{k}\|_{2}^{2}$ into the ridge regression framework without altering the underlying algorithm. By scaling the feature matrix and coefficients appropriately, we maintain the efficiency of the standard ridge regression solver while addressing the modified regularization objective.